# Cohet: A CXL-Driven Coherent Heterogeneous Computing Framework with Hardware-Calibrated Full-System Simulation

### v1

@@ I Introduction
With the end of Dennard scaling and slowdown of Moore’s law, homogeneous CPU architectures can no longer satisfy the explosive compute and memory demands of emerging data-driven workloads such as AI/ML and big data analytics [intel_accel_ecosystem]. This gap has spurred widespread adoption of heterogeneous systems that integrate CPUs with various XPUs (e.g., GPUs, SmartNICs/DPUs, and FPGAs) over high-speed interconnects such as PCIe to exploit their domain-specific accelerating capabilities [tpuv4, dpu, ipu, apunet, dnn_processor, geeps_dl, azure_accelerated], thus boosting overall computing performance. However, the conventional PCIe-based heterogeneous architecture suffers from two major limitations [demystifying_type2, mozart, schuh_cc-nic_2024, intel_accel_ecosystem, introduction_cxl, cuda_um, gh100_evaluation]. First, PCIe favors occasional host-device transfers of bulky data sets, making it well suited to bandwidth-intensive coarse-grained workloads. In contrast, latency-sensitive applications that require frequent or fine-grained CPU-XPU interactions are severely hindered by PCIe’s high latency and poor bandwidth in small data transfers. Second, PCIe is a noncoherent interconnect that has to rely on complex software stacks and significant programmer efforts to explicitly or implicitly manage data movement between host and accelerator, leading to a cumbersome programming model and poor development productivity.
@@
In recent years, both proprietary [nvlink_c2c, infinity_fabric, intel_qpi_upi, intel_accel_ecosystem] and open [CXL_spec, opencapi, genz, ccix] coherent interconnect technologies have emerged to address the limitations of PCIe. Proprietary solutions (e.g., NVIDIA’s NVLink-C2C [nvlink_c2c], AMD’s Infinity Fabric [infinity_fabric], and Intel’s UPI [intel_qpi_upi] and CMI [intel_accel_ecosystem]) deliver impressive performance, but lack hardware flexibility, software ecosystem diversity, and user code portability. Among open standards, CXL [CXL_spec] has become the de facto industry standard with over 240 active consortium members [cxl_integrators] after consolidating competing proposals OpenCAPI [opencapi], Gen-Z [genz], and CCIX [ccix]. Building upon the PCIe physical layer, CXL defines three sub-protocols: CXL.io, CXL.cache, and CXL.mem. To date, the majority of academic and industrial research works are focused on CXL-based disaggregated memory systems merely using CXL.mem [demystifying_type3, exploring_cxl_asic, samsung_cxl_ssd, maruf2023tpp, pond, lia_cxl_llm], which promise clear benefits in decoupling memory from compute and enabling independent scaling and sharing of memory resources. However, we contend that there is tremendous potential in exploiting CXL.cache together with CXL.mem to realize fully coherent heterogeneous computing. Such an approach can overcome the bottlenecks of the conventional heterogeneous architecture and fundamentally reshape the future cooperative computing paradigm.
@@
Despite its promise, CXL-based coherent heterogeneous computing faces three key challenges. First, the CXL hardware and software ecosystem remains immature. On the hardware side, CXL memory expanders have only reached early commercial availability [samsung_cmm_d, rambus, XConn, skhynix, Montage], and CXL-supported accelerators remain largely confined to FPGA platforms [agilex, cxl_fpga_ip] in experimental testbeds. On the software side, the OS kernel lacks native support for CXL accelerators [cxl_linux], and higher-level heterogeneous programming frameworks must adapt to CXL cache coherence and memory semantics [opencl, oneapi]. Second, academic researchers suffer from a dearth of realistic CXL platforms. Previous works often used a remote NUMA node as an emulator of CXL accelerators [huang_hal_2024, tarot_row_hammer, schuh_cc-nic_2024], which can introduce inaccuracies and misleading conclusions in some cases [demystifying_type2]. Finally, owing to the lack of effective evaluation and comprehensive analysis of CXL accelerators, the specific application scenarios benefiting from coherent heterogeneous computing remain unexplored.
@@
To address these challenges, this paper presents Cohet, the first CXL-driven coherent heterogeneous computing framework that is intended to deliver a software-hardware co-design solution targeting the open-source community. By interconnecting hosts and accelerators over CXL with cache-coherence, Cohet unifies their physical address spaces into a coherent memory pool. Cohet also provides a single shared memory view and a unified per-process page table to both CPUs and XPUs. The OS recognizes CPUs and XPUs as separate NUMA nodes, while providing standard malloc/mmap interfaces to programmers; memory allocation, page migration, and coherence management remain transparent. Moreover, we have developed SimCXL, an open-source full-system cycle-level simulator extended from gem5 that models all three CXL sub-protocols and three types of CXL devices. SimCXL has been rigorously calibrated against a real CXL-supported testbed with Intel FPGA accelerators and various memory expanders, achieving an average simulation error of 3%. Our evaluation shows that at cacheline granularity, CXL.cache reduces latency by 68% and increases bandwidth by 14.4 compared to DMA. To demonstrate the benefits of Cohet, we present a CXL-NIC design with two acceleration case-studies on SimCXL: 1) remote atomic operation (RAO), a key building block for distributed parallel computing [circustent, RAE, rao_sequencer, think_more_rdma, rdma_design, lock_2, ao_evaluating], and 2) remote procedure call (RPC), a fundamental inter-process communication layer in modern cloud microservices [micro_server, rpcnic, dagger, google_2023, protoAcc, meta_2020]. Our experimental results suggest that CXL-based RAO outperforms its PCIe-based counterpart by 5.5 to 40.2 across different operation types, and that CXL-based RPC achieves an average 1.86 speedup in (de)serialization compared to PCIe-based RPC.
@@ II-A1 CPU-XPU interactive interface
In PCIe-based systems, host and device memories reside in separate and non-coherent address spaces, with interactions realized by MMIO and DMA. MMIO exposes the base address registers (BARs) of a peripheral XPU device in the host address space, enabling the CPU to launch uncached loads and stores to the device BAR space. However, PCIe’s high per-transaction latency and strict write-ordering, which allows only one outstanding write, limit the MMIO performance [intel_accel_ecosystem]. On the XPU side, its DMA engine allows to directly access the host memory and delivers a much higher throughput for bulky contiguous transfers. But, DMA operations incur substantial per-transfer setup overhead, making fine-grained and randomly distributed data movements highly inefficient [demystifying_type2, understanding_pcie].
@@ II-B Compute Express Link
This paper focuses on type-1/2 devices, both of which can be recognized as accelerators (XPUs). Each device includes a host memory cache (HMC) that caches host memory and acts as a peer to the CPU’s L2 cache, with the CPU’s last-level cache (LLC) serving as a coherence synchronization point. The host tracks coherence among peer caches, while the CXL device interacts with the host over the device coherency engine (DCOH) using a lightweight MESI protocol. In addition to standard load/store operations, CXL.cache introduces the non-cacheable push (NC-P) instruction [cxl_fpga_ip], which allows a device to push a cacheline directly into the host LLC and invalidate it in the HMC, providing a low-latency path for returning computed results. Sections V-B will demonstrate how NC-P optimizes acceleration to boost application performance.
@@ III-A Limitations of Conventional Heterogeneous Computing
L1: Compute-memory coupling leads to resource under-utilization. Current compute units, regardless of CPUs or XPUs, are tightly coupled to their own memory, as illustrated with the dotted boxes in Fig. 1. These host servers and accelerators form basic performance scaling units, leading to either compute or memory over-provisions in today’s hyperscale datacenters [borg, pond]. In addition, the distributed physical memories remain mutually isolated, resulting in mandatory data copies and infeasibility of direct sharing of complex data structures in heterogeneous computing.
@@
L4: Complex programming model. The lack of hardware coherence results in complex programming models that depend on explicit/implicit memory management by programmers/software, further constrained by limited device memory capacity. Developers face a fundamental dilemma: manual data placement might yield high performance but impose significant programming efforts, whereas relying on software-managed unified memory mechanisms (e.g., CUDA unified memory (UM) [cuda_um]) simplifies development at the risk of performance loss due to expensive page fault handling.
@@ III-C1 Hardware level
One or more CXL switches compose a CXL fabric. A distributed resource scheduler (fabric manager) is implemented in each switch to allocate/release fabric-attached memory and XPU resources to a specific host. For each host (OS point of view), its CPUs and XPUs interconnect via the CXL fabric to form a unified physically-addressed heterogeneous memory pool, presenting a single memory view to all compute units. CXL’s hardware-maintained coherence not only allows CPUs and XPUs to cache and access each other’s memory at cacheline granularity but also supports cross-device atomic operations, enabling fine-grained and scalable synchronization across all running threads in the system. In this CXL-enabled architecture, the address translation service (ATS) lets CPUs and XPUs share a single per-process page table (denoted as Unified Page Table in the figure). When an XPU thread accesses a virtual address, it first looks up the mapping in its device-side address translation cache (ATC), analogous to the host TLB. Upon an ATC miss, the request is forwarded to the CPU-side IOMMU, which performs a page-table walk to resolve the physical address. The resolved mapping entry is then returned to the ATC, after which the XPU proceeds with its memory accesses [CXL_spec]. At this level, tailoring the microarchitecture of endpoint devices (e.g., XPUs and memory expanders) for target applications is also crucial for maximizing the benefits of the coherent interconnect, as demonstrated with our CXL-NIC designs for accelerating two killer-apps in Section V.
@@ III-C2 OS level
The Linux kernel recognizes CPUs and XPUs as separate NUMA nodes. Heterogeneous memory management (HMM) merges device memory with host memory into the system memory pool, maintains the unified page table, and exposes a standard mmap/malloc interface to upper-layer software [linux_hmm]. Specifically, the device driver probes the device memory size during system initialization, registers a device instance with HMM, and implements the callback functions required for unified page-table management, such as handling ATC invalidations and page migrations. Furthermore, we modify the kernel’s numa_init routine, which inspects the available system memory, initializes the host and device memory as distinct NUMA nodes based on their types, and binds them to the corresponding CPU or XPU.
A malloc call allocates a page-table entry without assigning a physical frame, allowing memory overcommitment. On an XPU’s first access to a given virtual address, an ATC miss triggers an IOMMU translation request. The kernel then updates the page-table entry to point to XPU physical memory. Once pages are allocated, both CPUs and XPUs access them at cacheline granularity under CXL’s coherence guarantee. By contrast, PCIe-based non-coherent interconnects incur high-overhead OS page-fault interrupts and DMA-based page copies, which prior studies identify as the primary performance bottleneck in frameworks such as UM [um_overhead]. When the unified page table is about to be updated due to page migration or swapping, HMM invokes the registered driver callback. The driver then temporarily blocks the device from accessing the affected page-table entries, allowing HMM to safely perform the update and trigger the IOMMU invalidation process. According to the ATS protocol, this process invalidates the corresponding entries in the device-side ATC. Once the invalidation has been completed, HMM notifies the driver to resume device address translation. Moreover, adaptive page migration in Cohet is a potential performance optimization, left for future work.
@@ III-C3 User level
Cohet embraces OpenCL 3.0 [opencl], the most pervasive cross-vendor open standard for low-level heterogeneous parallel programming, to deliver a fully transparent programming model. While the standard OpenCL requires special APIs for explicit (such as clHostMemAllocINTEL) or implicit memory management (such as clSharedMemAllocINTEL) before execution, Cohet allows applications to use mmap, malloc, and C++ new exactly as in CPU-only systems, without incurring page-fault penalties. Fig. 4 compares Cohet with CUDA’s explicit and implicit copy programming models, using AXPY operation () as a pseudocode example. Fig. 4(a) relies heavily on programmers for manual memory management, resulting in the highest code complexity (16 lines). Fig. 4(b) employs UM to reduce programming effort (10 lines), but it incurs non-negligible overhead from implicit data copies. In contrast, Cohet in Fig. 4(c) further simplifies heterogeneous programming to a level approaching homogeneous programming (9 lines), while achieving performance comparable to manually managed data.
@@ IV-A Simulator Architecture
SimCXL is a cycle-level simulator intended for modeling the proposed Cohet computing framework with a CXL interconnect fabric connecting CPUs, XPUs, and a memory pool in a cache-coherent manner. Fig. 5 illustrates its modular architecture built on gem5’s full-system (FS) mode [gem5_paper, gem5_v20], modeling both the guest OS and hardware architecture to enable realistic simulation of software-hardware interactions.
@@ IV-B CXL Protocol Supports
Fig. 6 shows an x86 platform with a Ruby subsystem used in the FS mode. Ruby [gem5_ruby] supports the flexible modeling of various cache coherence protocols by specifying state transitions of multi-level caches using the domain-specific language SLICC. On the host side, the CPU cores and their MMUs connect to the Ruby sequencer through cache and Walker ports, respectively. The sequencer routes memory access requests by analyzing packet destination addresses: memory access requests are converted into RubyRequest packets and forwarded to the cache hierarchy (red arrows), while IO accesses are directed to target devices via the IO bus (orange arrows). The memory interface coordinates requests from LLC and DMA controllers while converting RubyRequest packets into gem5 memory packets for memory controllers. The device side includes the CXL accelerator with its HMC and device memory. The accelerator utilizes three CXL sub-protocols. CXL.io supports MMIO and DMA through PIO and DMA ports, respectively. CXL.cache enables coherent memory access in cacheline granularity via the cache port. CXL.mem incorporates device memory into the host memory address space, managed uniformly by the memory interface, allowing cores to perform standard load/store operations with full cache coherence.

### v2

@@ I Introduction
With the end of Dennard scaling and slowdown of Moore’s law, homogeneous CPU architectures can no longer satisfy the explosive compute and memory demands of emerging data-driven workloads such as AI/ML and big data analytics [95]. This gap has spurred widespread adoption of heterogeneous systems that integrate CPUs with various XPUs (e.g., GPUs, SmartNICs/DPUs, and FPGAs) over high-speed interconnects such as PCIe to exploit their domain-specific accelerating capabilities [37, 8, 7, 29, 24, 17, 22], thus boosting overall computing performance. However, the conventional PCIe-based heterogeneous architecture suffers from two major limitations [36, 79, 67, 95, 70, 58, 66]. First, PCIe favors occasional host-device transfers of bulky data sets, making it well suited to bandwidth-intensive coarse-grained workloads. In contrast, latency-sensitive applications that require frequent or fine-grained CPU-XPU interactions are severely hindered by PCIe’s high latency and poor bandwidth in small data transfers. Second, PCIe is a noncoherent interconnect that has to rely on complex software stacks and significant programmer efforts to explicitly or implicitly manage data movement between host and accelerator, leading to a cumbersome programming model and poor development productivity.
@@
In recent years, both proprietary [57, 4, 31, 95] and open [18, 60, 28, 10] coherent interconnect technologies have emerged to address the limitations of PCIe. Proprietary solutions (e.g., NVIDIA’s NVLink-C2C [57], AMD’s Infinity Fabric [4], and Intel’s UPI [31] and CMI [95]) deliver impressive performance, but lack hardware flexibility, software ecosystem diversity, and user code portability. Among open standards, CXL [18] has become the de facto industry standard with over 240 active consortium members [34] after consolidating competing proposals OpenCAPI [60], Gen-Z [28], and CCIX [10]. Building upon the PCIe physical layer, CXL defines three sub-protocols: CXL.io, CXL.cache, and CXL.mem. To date, the majority of academic and industrial research works are focused on CXL-based disaggregated memory systems merely using CXL.mem [78, 81, 96, 51, 44, 41], which promise clear benefits in decoupling memory from compute and enabling independent scaling and sharing of memory resources. However, we contend that there is tremendous potential in exploiting CXL.cache together with CXL.mem to realize fully coherent heterogeneous computing. Such an approach can overcome the bottlenecks of the conventional heterogeneous architecture and fundamentally reshape the future cooperative computing paradigm.
@@
Despite its promise, CXL-based coherent heterogeneous computing faces three key challenges. First, the CXL hardware and software ecosystem remains immature. On the hardware side, CXL memory expanders have only reached early commercial availability [14, 19, 89, 72, 52], and CXL-supported accelerators remain largely confined to FPGA platforms [2, 15] in experimental testbeds. On the software side, the OS kernel lacks native support for CXL accelerators [16], and higher-level heterogeneous programming frameworks must adapt to CXL cache coherence and memory semantics [61, 59]. Second, academic researchers suffer from a dearth of realistic CXL platforms. Previous works often used a remote NUMA node as an emulator of CXL accelerators [33, 74, 67], which can introduce inaccuracies and misleading conclusions in some cases [36]. Finally, owing to the lack of effective evaluation and comprehensive analysis of CXL accelerators, the specific application scenarios benefiting from coherent heterogeneous computing remain unexplored.
@@
To address these challenges, this paper presents Cohet, the first CXL-driven coherent heterogeneous computing framework that is intended to deliver a software-hardware co-design solution targeting the open-source community. By interconnecting hosts and accelerators over CXL with cache-coherence, Cohet unifies their physical address spaces into a coherent memory pool. Cohet also provides a single shared memory view and a unified per-process page table to both CPUs and XPUs. The OS recognizes CPUs and XPUs as separate NUMA nodes, while providing standard malloc/mmap interfaces to programmers; memory allocation, page migration, and coherence management remain transparent. Moreover, we have developed SimCXL, an open-source full-system cycle-level simulator extended from gem5 that models all three CXL sub-protocols and three types of CXL devices. SimCXL has been rigorously calibrated against a real CXL-supported testbed with Intel FPGA accelerators and various memory expanders, achieving an average simulation error of 3%. Our evaluation shows that at cacheline granularity, CXL.cache reduces latency by 68% and increases bandwidth by 14.4 compared to DMA. To demonstrate the benefits of Cohet, we present a CXL-NIC design with two acceleration case-studies on SimCXL: 1) remote atomic operation (RAO), a key building block for distributed parallel computing [88, 85, 11, 49, 38, 56, 68], and 2) remote procedure call (RPC), a fundamental inter-process communication layer in modern cloud microservices [62, 98, 42, 69, 39, 76]. Our experimental results suggest that CXL-based RAO outperforms its PCIe-based counterpart by 5.5 to 40.2 across different operation types, and that CXL-based RPC achieves an average 1.86 speedup in (de)serialization compared to PCIe-based RPC.
@@ II-A1 CPU-XPU interactive interface
In PCIe-based systems, host and device memories reside in separate and non-coherent address spaces, with interactions realized by MMIO and DMA. MMIO exposes the base address registers (BARs) of a peripheral XPU device in the host address space, enabling the CPU to launch uncached loads and stores to the device BAR space. However, PCIe’s high per-transaction latency and strict write-ordering, which allows only one outstanding write, limit the MMIO performance [95]. On the XPU side, its DMA engine allows to directly access the host memory and delivers a much higher throughput for bulky contiguous transfers. But, DMA operations incur substantial per-transfer setup overhead, making fine-grained and randomly distributed data movements highly inefficient [36, 32].
@@ II-B Compute Express Link
This paper focuses on type-1/2 devices, both of which can be recognized as accelerators (XPUs). Each device includes a host memory cache (HMC) that caches host memory and acts as a peer to the CPU’s L2 cache, with the CPU’s last-level cache (LLC) serving as a coherence synchronization point. The host tracks coherence among peer caches, while the CXL device interacts with the host over the device coherency engine (DCOH) using a lightweight MESI protocol. In addition to standard load/store operations, CXL.cache introduces the non-cacheable push (NC-P) instruction [15], which allows a device to push a cacheline directly into the host LLC and invalidate it in the HMC, providing a low-latency path for returning computed results. Sections V-B will demonstrate how NC-P optimizes acceleration to boost application performance.
@@ III-A Limitations of Conventional Heterogeneous Computing
L1: Compute-memory coupling leads to resource under-utilization. Current compute units, regardless of CPUs or XPUs, are tightly coupled to their own memory, as illustrated with the dotted boxes in Fig. 1. These host servers and accelerators form basic performance scaling units, leading to either compute or memory over-provisions in today’s hyperscale datacenters [82, 44]. In addition, the distributed physical memories remain mutually isolated, resulting in mandatory data copies and infeasibility of direct sharing of complex data structures in heterogeneous computing.
@@
L4: Complex programming model. The lack of hardware coherence results in complex programming models that depend on explicit/implicit memory management by programmers/software, further constrained by limited device memory capacity. Developers face a fundamental dilemma: manual data placement might yield high performance but impose significant programming efforts, whereas relying on software-managed unified memory mechanisms (e.g., CUDA unified memory (UM) [58]) simplifies development at the risk of performance loss due to expensive page fault handling.
@@ III-C1 Hardware level
One or more CXL switches compose a CXL fabric. A distributed resource scheduler (fabric manager) is implemented in each switch to allocate/release fabric-attached memory and XPU resources to a specific host. For each host (OS point of view), its CPUs and XPUs interconnect via the CXL fabric to form a unified physically-addressed heterogeneous memory pool, presenting a single memory view to all compute units. CXL’s hardware-maintained coherence not only allows CPUs and XPUs to cache and access each other’s memory at cacheline granularity but also supports cross-device atomic operations, enabling fine-grained and scalable synchronization across all running threads in the system. In this CXL-enabled architecture, the address translation service (ATS) lets CPUs and XPUs share a single per-process page table (denoted as Unified Page Table in the figure). When an XPU thread accesses a virtual address, it first looks up the mapping in its device-side address translation cache (ATC), analogous to the host TLB. Upon an ATC miss, the request is forwarded to the CPU-side IOMMU, which performs a page-table walk to resolve the physical address. The resolved mapping entry is then returned to the ATC, after which the XPU proceeds with its memory accesses [18]. At this level, tailoring the microarchitecture of endpoint devices (e.g., XPUs and memory expanders) for target applications is also crucial for maximizing the benefits of the coherent interconnect, as demonstrated with our CXL-NIC designs for accelerating two killer-apps in Section V.
@@ III-C2 OS level
The Linux kernel recognizes CPUs and XPUs as separate NUMA nodes. Heterogeneous memory management (HMM) merges device memory with host memory into the system memory pool, maintains the unified page table, and exposes a standard mmap/malloc interface to upper-layer software [45]. Specifically, the device driver probes the device memory size during system initialization, registers a device instance with HMM, and implements the callback functions required for unified page-table management, such as handling ATC invalidations and page migrations. Furthermore, we modify the kernel’s numa_init routine, which inspects the available system memory, initializes the host and device memory as distinct NUMA nodes based on their types, and binds them to the corresponding CPU or XPU.
A malloc call allocates a page-table entry without assigning a physical frame, allowing memory overcommitment. On an XPU’s first access to a given virtual address, an ATC miss triggers an IOMMU translation request. The kernel then updates the page-table entry to point to XPU physical memory. Once pages are allocated, both CPUs and XPUs access them at cacheline granularity under CXL’s coherence guarantee. By contrast, PCIe-based non-coherent interconnects incur high-overhead OS page-fault interrupts and DMA-based page copies, which prior studies identify as the primary performance bottleneck in frameworks such as UM [40]. When the unified page table is about to be updated due to page migration or swapping, HMM invokes the registered driver callback. The driver then temporarily blocks the device from accessing the affected page-table entries, allowing HMM to safely perform the update and trigger the IOMMU invalidation process. According to the ATS protocol, this process invalidates the corresponding entries in the device-side ATC. Once the invalidation has been completed, HMM notifies the driver to resume device address translation. Moreover, adaptive page migration in Cohet is a potential performance optimization, left for future work.
@@ III-C3 User level
Cohet embraces OpenCL 3.0 [61], the most pervasive cross-vendor open standard for low-level heterogeneous parallel programming, to deliver a fully transparent programming model. While the standard OpenCL requires special APIs for explicit (such as clHostMemAllocINTEL) or implicit memory management (such as clSharedMemAllocINTEL) before execution, Cohet allows applications to use mmap, malloc, and C++ new exactly as in CPU-only systems, without incurring page-fault penalties. Fig. 4 compares Cohet with CUDA’s explicit and implicit copy programming models, using AXPY operation () as a pseudocode example. Fig. 4(a) relies heavily on programmers for manual memory management, resulting in the highest code complexity (16 lines). Fig. 4(b) employs UM to reduce programming effort (10 lines), but it incurs non-negligible overhead from implicit data copies. In contrast, Cohet in Fig. 4(c) further simplifies heterogeneous programming to a level approaching homogeneous programming (9 lines), while achieving performance comparable to manually managed data.
@@ IV-A Simulator Architecture
SimCXL is a cycle-level simulator intended for modeling the proposed Cohet computing framework with a CXL interconnect fabric connecting CPUs, XPUs, and a memory pool in a cache-coherent manner. Fig. 5 illustrates its modular architecture built on gem5’s full-system (FS) mode [6, 48], modeling both the guest OS and hardware architecture to enable realistic simulation of software-hardware interactions.
@@ IV-B CXL Protocol Supports
Fig. 6 shows an x86 platform with a Ruby subsystem used in the FS mode. Ruby [27] supports the flexible modeling of various cache coherence protocols by specifying state transitions of multi-level caches using the domain-specific language SLICC. On the host side, the CPU cores and their MMUs connect to the Ruby sequencer through cache and Walker ports, respectively. The sequencer routes memory access requests by analyzing packet destination addresses: memory access requests are converted into RubyRequest packets and forwarded to the cache hierarchy (red arrows), while IO accesses are directed to target devices via the IO bus (orange arrows). The memory interface coordinates requests from LLC and DMA controllers while converting RubyRequest packets into gem5 memory packets for memory controllers. The device side includes the CXL accelerator with its HMC and device memory. The accelerator utilizes three CXL sub-protocols. CXL.io supports MMIO and DMA through PIO and DMA ports, respectively. CXL.cache enables coherent memory access in cacheline granularity via the cache port. CXL.mem incorporates device memory into the host memory address space, managed uniformly by the memory interface, allowing cores to perform standard load/store operations with full cache coherence.